OpenStudy (anonymous):

x^2-5kx+25 is the square of binomial. What is a possible value of k ? Please include steps, thanks!

OpenStudy (anonymous):

Its\[(x-5)^2 = x^2+25-2(x)(5) = x^2+25-10x so k =2

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[(a-b)^2 = a^2+b^2-2ab . This is a formula.

OpenStudy (anonymous):

what formula?

OpenStudy (anonymous):

(a-b)^2 = a^2+b^2-2ab

OpenStudy (nikvist):

\[k=\pm 2\]